# D-SPEAR: Dual-Stream Prioritized Experience Adaptive Replay for Stable Reinforcement Learning in Robotic Manipulation

### 28 Mar 2026 v1

@@ II-B Dual-Stream Experience Replay
with size $\lfloor\lambda N\rfloor$, where $N$ denotes the batch size and $\lambda\in[0,1]$ controls the proportion of uniformly sampled transitions. In practice, we construct a candidate set whose size is four times the batch size (candidate sampling ratio = 4) before applying the prioritized sampling strategies.
@@ II-D Robust Critic Objective
0: Batch size $N$, minimum anchor ratio $\lambda_{\min}$, priority   exponents $\alpha_{c}$, $\beta_{a}$, Huber threshold $\delta$
@@ III-A Experimental Setup
All environments are controlled using an operational space controller in position mode (OSC_POSITION). The control frequency is set to $20$ Hz, and each episode has a fixed horizon of $500$ time steps. Visual observations are disabled, and only low-dimensional proprioceptive and object-state observations are used. Reward shaping provided by robosuite is enabled to facilitate learning while preserving task difficulty. A summary of the environment configuration is provided in Table I.
@@
D-SPEAR is implemented on top of a Soft Actor–Critic (SAC) backbone. Both the actor and critic networks are parameterized as two-layer multilayer perceptrons with $256$ hidden units per layer. A shared replay buffer stores transitions of the form $(s_{t},a_{t},r_{t},s_{t+1},d_{t})$. The proposed dual-stream replay mechanism constructs separate mini-batches for actor and critic updates, as described in Section III. All algorithmic and training hyperparameters are summarized in Table II.
@@
All methods are trained for $500$ episodes with a maximum of $500$ steps per episode, resulting in a total interaction budget of $2.5\times 10^{5}$ environment steps. The first 5,0005{,}000 steps are used for random exploration before policy updates begin. After the warm-up phase, one gradient update is performed per environment step. Each experiment is repeated over $5$ random seeds. Performance is evaluated using episode return, reported as the mean across seeds.
@@ III-B Main Results on Robotic Manipulation Tasks
Figure 3: Performance comparison on robosuite manipulation tasks. Episode return as a function of environment steps on Lift and Door. Solid lines denote the mean performance over $5$ random seeds, and shaded regions indicate one standard deviation.
@@ III-C Ablation Study
Figure 4: Ablation study on robosuite manipulation tasks. We compare the full D-SPEAR framework with variants that remove key components: w/o dual-stream, w/o low-actor, and w/o high-critic. Results are averaged over $5$ random seeds, with shaded regions indicating one standard deviation.

### 01 Apr 2026 v2

@@ II-B Dual-Stream Experience Replay
with size $\lfloor\lambda N\rfloor$, where NN denotes the batch size and $\lambda\in[0,1]$ controls the proportion of uniformly sampled transitions. In practice, we construct a candidate set whose size is four times the batch size (candidate sampling ratio = 4) before applying the prioritized sampling strategies.
@@ II-D Robust Critic Objective
0: Batch size NN, minimum anchor ratio $\lambda_{\min}$, priority   exponents $\alpha_{c}$, $\beta_{a}$, Huber threshold $\delta$
@@ III-A Experimental Setup
All environments are controlled using an operational space controller in position mode (OSC_POSITION). The control frequency is set to 2020 Hz, and each episode has a fixed horizon of 500500 time steps. Visual observations are disabled, and only low-dimensional proprioceptive and object-state observations are used. Reward shaping provided by robosuite is enabled to facilitate learning while preserving task difficulty. A summary of the environment configuration is provided in Table I.
@@
D-SPEAR is implemented on top of a Soft Actor–Critic (SAC) backbone. Both the actor and critic networks are parameterized as two-layer multilayer perceptrons with 256256 hidden units per layer. A shared replay buffer stores transitions of the form $(s_{t},a_{t},r_{t},s_{t+1},d_{t})$. The proposed dual-stream replay mechanism constructs separate mini-batches for actor and critic updates, as described in Section III. All algorithmic and training hyperparameters are summarized in Table II.
@@
All methods are trained for 500500 episodes with a maximum of 500500 steps per episode, resulting in a total interaction budget of $2.5\times 10^{5}$ environment steps. The first 5,0005{,}000 steps are used for random exploration before policy updates begin. After the warm-up phase, one gradient update is performed per environment step. Each experiment is repeated over 55 random seeds. Performance is evaluated using episode return, reported as the mean across seeds.
@@ III-B Main Results on Robotic Manipulation Tasks
Figure 3: Performance comparison on robosuite manipulation tasks. Episode return as a function of environment steps on Lift and Door. Solid lines denote the mean performance over 55 random seeds, and shaded regions indicate one standard deviation.
@@ III-C Ablation Study
Figure 4: Ablation study on robosuite manipulation tasks. We compare the full D-SPEAR framework with variants that remove key components: w/o dual-stream, w/o low-actor, and w/o high-critic. Results are averaged over 55 random seeds, with shaded regions indicating one standard deviation.